Consider the country of Agrovia, whose GDP is discussed in “A Numerical Example” on page 396. Construct a set of national accounts like that in Table 20-6 assuming that wheat costs $5 per bushel, there is no depreciation, wages are three-fourths of national output, indirect business taxes are used to finance 100 percent of government spending, and the balance of income goes as rent income to farmers.
1
Expert's answer
2021-09-13T18:41:21-0400

Output = $1000

Cost of FP = $750 in wages, $125 in interest, and $75 in rent

"Profit = output \u2013 expenses = $ 1000 \u2013 ($ 750 + $ 125 + $ 75) = $ 50"

Investment & Consumption

"C=\u00bex1000=750"

"I=1000\u2013750=250"

Product Approaches  the GDP with Product Approaches is 1000

"Income Approaches GDP\/Y = w + r + I + pGDP\/Y=w+r+I+p"

"GDP\/Y=750+125+75+50=1000"